# Low Regularity of Self-Similar Solutions of Two-Dimensional Riemann problems with Shocks for the Isentropic Euler system

### v1

@@ 4.2. Lower Regularity of the Prantl-Meyer Reflection Solutions for Supersonic Flows past a Solid Ramp
The second example is the Prandtl reflection problem for the isentropic Euler system (2.1). This is of a self-similar structure that occurs when a $2$-D supersonic flow with density $\rho_{\infty}>0$ and velocity ${\mathbf{v}}_{\infty}=(u_{\infty},0)$, $u_{\infty}>0$, along the wedge-axis hits the wedge in the direction at $t=0$. See Figs. 4.64.6; also see Bae-Chen-Feldman [2, 3] and Elling-Liu [26].
Figure 4.5. Prandtl supersonic shock reflection
Figure 4.6. Prandtl subsonic shock reflection
@@
First, a similar argument as made for the regular shock reflection at the reflection point $P_{0}$ (see Figs. 4.24.2) yields that, at wedge-vertex $P_{1}$, for a given uniform incoming flow $(\rho_{\infty},{\mathbf{v}}_{\infty})$, there is a detachment angle $\theta_{\rm w}^{\rm d}\in(0,\frac{\pi}{2})$ such that the system of algebraic equations (2.9) and (2.13) for state $(O)$ has two solutions for each wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$ such that the entropy condition is satisfied for the two-shock configuration. The weak state $(O)$ with the smaller density is expected to be physical. So we always refer to state $(O)$ as the weak state $(O)$. Similarly, it follows from the Rankine-Hugoniot conditions (2.9) and the slip boundary condition (2.13) that
The shock line $\Gamma_{\rm shock}^{O}$ between states $(O)$ and ()($\infty$)( ∞ ) is not parallel to $\Gamma_{\rm wedge}$ for all $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$. (4.5)
Depending on the wedge-angle, state $(O)$ can be either supersonic or subsonic at $P_{1}$. It determines the type of the reflection, supersonic or subsonic, as shown in Fig. 4.6 and Fig. 4.6, respectively.
Second, by a straightforward computation, we know that there exists a unique constant state $(N)$, which determines the normal reflection of state ()($\infty$)( ∞ ) from the wedge boundary $\Gamma_{\rm wedge}$ so that state $(N)$ satisfies the slip boundary condition (2.13) along $\Gamma_{\rm wedge}$ and the Rankine-Hugoniot conditions (2.9) along a straight line $\Gamma_{\rm shock}^{N}$, which lies in $\Lambda$ and is parallel to $\Gamma_{\rm wedge}$.
@@
    If state $(O)$ at $P_{1}$ is supersonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|>c_{O}$, the solution has the supersonic reflection structure as in Fig. 4.6 at point $P_{1}$. If state $(O)$ at $P_{1}$ is subsonic or sonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|\leq c_{O}$, the solution has the subsonic reflection structure as in Fig. 4.6.
@@
    The solution coincides with states ()($\infty$)( ∞ ), $(O)$, and $(N)$ in their respective regions. Specifically,
@@
    and
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{O},{\mathbf{v}}_{O})(P_{1})\qquad\mbox{if state (O) at $P_{1}$ is subsonic or sonic}.$
@@
We remark that, for the supersonic reflections, it follows from (ii)–(iv) of the definition above and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arcs $\Gamma_{\rm sonic}^{N}$ and $\Gamma_{\rm sonic}^{O}$.
@@
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
@@
Then we can follow the proof of Theorem 4.10 to show that the regular solutions satisfy the assumptions of Theorem 3.2 similarly. The only (slight) difference is in showing that condition (b) of Lemma 4.1 holds. We need to show that $\Gamma_{\rm shock}$ is not a straight segment. Let us first consider the supersonic Prandtl reflection case; see Fig. 4.6. Arguing as in the proof of Theorem 4.10, we show that, at point $P_{4}$ (resp. $P_{5}$), curve $\Gamma_{\rm shock}$ is tangential to line $\Gamma_{\rm shock}^{N}$ (resp. $\Gamma_{\rm shock}^{O}$). If $\Gamma_{\rm shock}$ is a straight segment, we obtain that $\Gamma_{\rm shock}^{N}$ and $\Gamma_{\rm shock}^{O}$ lie in the same line. However, $\Gamma_{\rm shock}^{N}$ is parallel to $\Gamma_{\rm wedge}$ with a positive distance from it, while $\Gamma_{\rm shock}^{O}$ passes through point $P_{1}\in\Gamma_{\rm wedge}$. This contradiction shows that $\Gamma_{\rm shock}$ is not a straight segment. In the case of subsonic or sonic Prandtl reflection as in Fig. 4.6, we argue similarly, except we consider point $P_{1}$ instead of point $P_{5}$ and use the fact that, in the subsonic or sonic case, $\Gamma_{\rm shock}$ is parallel to $\Gamma_{\rm shock}^{O}$ at $P_{1}$ by the last equality in condition (iv) of the definition of admissible solutions. Then we conclude in the same way as in the case of supersonic Prandtl reflection. Thus, we do not repeat the similar arguments for the proof. ∎
@@ 4.3. Lower Regularity of the Shock Diffraction Solutions of the Lighthill Problem
The third problem is the Lighthill problem, i.e., the shock diffraction problem, for the isentropic Euler system. As discussed in [16] for the potential flow and shown in Fig. 4.7, the Lighthill shock diffraction problem arises as a straight incident shock passes through a wedge stepping down.
Figure 4.7. Lighthill shock diffraction problem
Initially, we consider two piecewise constant Riemann data with the left state $(1)$: $(\rho_{1},u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$ and the right state $(0)$: $(\rho_{0},0,0)$, separated by a vertical shock, which hits the wedge at the wedge-corner $P_{1}$. Then the Lighthill shock diffraction problem evolves in a self-similar structure as time goes on. In the self-similar coordinates ${\boldsymbol{\xi}}$, the incident shock $\Gamma_{\rm shock}^{1}$ is given by $\xi_{1}=\xi_{1}^{0}$. By a straightforward calculation, it follows from the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12) that the location of the incident shock satisfies $0<\xi_{1}^{0}<c_{1}$, where $c_{1}=\rho_{1}^{\frac{\gamma-1}{2}}$ is the sonic speed of state $(1)$. Thus, as shown in Fig. 4.7, the incident shock $\Gamma_{\rm shock}^{1}$ interacts with the sonic circle $\Gamma_{\rm sonic}$ of state $(1)$ and becomes a transonic shock $\Gamma_{\rm shock}$, and then $\Gamma_{\rm shock}$ meets wedge $\Gamma_{\rm wedge}^{0}$ perpendicularly.
Let $\Lambda_{0}$, $\Lambda_{1}$, and $\Omega$ be defined, as shown in Fig. 4.7. Let $\Lambda=\overline{\Lambda_{0}\cup\Lambda_{1}\cup\Omega}$. An entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in $\Lambda$ in the sense of Definition 2.2 is called a regular shock diffraction solution of the Lighthill problem for the isentropic Euler system if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
@@
    $\rho,{\mathbf{v}}\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})\setminus\{P_{3}\}\right)\cap C^{0,1}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})$ for some $\sigma>0$;
@@
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
@@ 4.4. Lower Regularity of the Riemann Solutions with Four-Shock Interactions
where $\theta$ is the polar angle of point ${\mathbf{x}}\in\mathbb{R}^{2}$ and the four parameters $\theta_{12},\theta_{32},\theta_{34},\theta_{14}\in(0,\frac{\pi}{2})$. On each $\Lambda_{i}$, suitable constant states $(i)$ with values $(\rho_{i},{\mathbf{v}}_{i})$ are given for the Riemann initial data such that any two neighboring states are connected by exactly one planar shock discontinuity, which satisfies the Rankine-Hugoniot conditions and the entropy condition for the isentropic Euler system. As introduced in [12], under the symmetry assumption that $\theta_{12}=\theta_{14}=\theta_{1}$, $\theta_{32}=\theta_{34}=\theta_{2}$, and states in $\Lambda_{2}$ and $\Lambda_{4}$ are the same constants, and under a structure assumption that one forward shock is generated between states $(1)$ and $(j)$ and one backward shock is generated between states $(3)$ and $(j)$, governed by the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12), we expect the Riemann problem develops a Riemann solution with the structure as shown in Fig. 4.8.
Figure 4.8. Riemann problem with four-shock interactions
On the symmetry line that is the dashed line shown in Fig. 4.8, the velocity satisfies (2.13) with $\boldsymbol{\nu}$ being the vertical direction. At the interaction points $P_{1}$ and $P_{4}$, based on an argument similar to the one for the regular shock reflection problem at the reflection point $P_{0}$, with the symmetry line and state $(2)$ in Fig. 4.8 corresponding to $\Gamma_{\rm wedge}$ and state $(1)$ in Fig. 4.2, there exists a detachment angle $\theta^{\rm d}\in(0,\frac{\pi}{2})$ depending on the data such that, for $\theta_{1},\theta_{2}\in(0,\theta_{\rm w}^{\rm d})$, there exist two constant states $(5)$ and $(6)$ in $\Lambda_{5}$ and $\Lambda_{6}$, respectively, satisfying the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12). One corresponds to the strong shock and the other to the weak shock. We always select the weak one because it expects to be stable in general. Then there exists a sonic angle $\theta^{\rm s}\in(0,\theta^{\rm d}]$ such that the state at $P_{1}$ (or $P_{4}$) is pseudo-supersonic if $\theta_{1}\in(0,\theta^{\rm s})$ (or $\theta_{2}\in(0,\theta^{\rm s})$) and is pseudo-subsonic if $\theta_{1}\in(\theta^{\rm s},\theta^{\rm d}]$ (or $\theta_{2}\in(\theta^{\rm s},\theta^{\rm d}]$).
@@
The tangents to Γshock1 at P2 and P3 are not parallel to each $\mbox{\it The tangents to $\Gamma_{\rm shock}^{1}$ at $P_{2}$ and $P_{3}$ are not parallel to each other},$ (4.6)
@@
Let $\Lambda_{i}$ for $i=1,\cdots,6$, if they exist, and let $\Omega$ be defined as shown in Fig. 4.8. Let $\Lambda=\mathbb{R}^{2}$. An entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in $\Lambda$ in the sense of Definition 2.2 is called a regular Riemann solution with four-shock interactions for the isentropic Euler system if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
@@
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{in}\,\,\,\Lambda_{1},\\[1.42262pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{2},\\[1.42262pt] (\rho_{3},{\mathbf{v}}_{3})\qquad\mbox{in}\,\,\,\Lambda_{3},\\[1.42262pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{4},\\[1.42262pt] (\rho_{5},{\mathbf{v}}_{5})\qquad\mbox{in}\,\,\,\Lambda_{5},\quad\mbox{if state (5) at $P_{4}$ is supersonic}\\[1.42262pt] (\rho_{6},{\mathbf{v}}_{6})\qquad\mbox{in}\,\,\,\Lambda_{6},\quad\mbox{if state (6) at $P_{1}$ is supersonic},\end{cases}$
    where
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{4}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{5},{\mathbf{v}}_{5})(P_{4})\qquad\mbox{if state (5) at $P_{4}$ is subsonic or sonic},$
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{6},{\mathbf{v}}_{6})(P_{1})\qquad\mbox{if state (6) at $P_{1}$ is subsonic or sonic}.$
@@
If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both supersonic, then $M=6$, ${N_{1}}=0$, and ${N_{2}}=4$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig. 4.8, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{6}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{4}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both subsonic, then $M=4$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,2,3,4$) being given as in Fig. 4.8, $\Lambda_{5}$ and $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{1}$.
If state $(5)$ at $P_{4}$ is supersonic and state $(6)$ at $P_{1}$ is subsonic, then $M=5$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,\cdots,5$) being given as in Fig. 4.8, $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{3}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{1}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
It state $(5)$ at $P_{4}$ is subsonic and state $(6)$ at $P_{1}$ is supersonic, then $M=5$, ${N_{1}}=0$ and ${N_{2}}=3$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig. 4.8, $\Lambda_{5}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$.
@@
    $(\rho,{\mathbf{v}})\in C^{1}\big{(}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})\setminus\{P_{2},P_{3},P_{5},P_{6}\}\big{)}\cap C^{0,1}\big{(}{\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega}\big{)}$ for some $\sigma>0$, where $\Gamma_{\rm sonic}^{i}=\emptyset$ when it does not exist for $i=1,2$;
@@
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.

### v2

@@ 4.2. Lower Regularity of the Prantl-Meyer Reflection Solutions for Supersonic Flows past a Solid Ramp
The second example is the Prandtl reflection problem for the isentropic Euler system (2.1). This is of a self-similar structure that occurs when a $2$-D supersonic flow with density $\rho_{\infty}>0$ and velocity ${\mathbf{v}}_{\infty}=(u_{\infty},0)$, $u_{\infty}>0$, along the wedge-axis hits the wedge in the direction at $t=0$. See Figs. 77; also see Bae-Chen-Feldman [2, 3] and Elling-Liu [26].
Figure 6. Prandtl supersonic shock reflection
Figure 7. Prandtl subsonic shock reflection
@@
First, a similar argument as made for the regular shock reflection at the reflection point $P_{0}$ (see Figs. 33) yields that, at wedge-vertex $P_{1}$, for a given uniform incoming flow $(\rho_{\infty},{\mathbf{v}}_{\infty})$, there is a detachment angle $\theta_{\rm w}^{\rm d}\in(0,\frac{\pi}{2})$ such that the system of algebraic equations (2.9) and (2.13) for state $(O)$ has two solutions for each wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$ such that the entropy condition is satisfied for the two-shock configuration. The weak state $(O)$ with the smaller density is expected to be physical. So we always refer to state $(O)$ as the weak state $(O)$. Similarly, it follows from the Rankine-Hugoniot conditions (2.9) and the slip boundary condition (2.13) that
The shock line $\Gamma_{\rm shock}^{O}$ between states $(O)$ and $(\infty)$ is not parallel to $\Gamma_{\rm wedge}$ for all $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$. (4.5)
Depending on the wedge-angle, state $(O)$ can be either supersonic or subsonic at $P_{1}$. It determines the type of the reflection, supersonic or subsonic, as shown in Fig. 7 and Fig. 7, respectively.
Second, by a straightforward computation, we know that there exists a unique constant state $(N)$, which determines the normal reflection of state $(\infty)$ from the wedge boundary $\Gamma_{\rm wedge}$ so that state $(N)$ satisfies the slip boundary condition (2.13) along $\Gamma_{\rm wedge}$ and the Rankine-Hugoniot conditions (2.9) along a straight line $\Gamma_{\rm shock}^{N}$, which lies in $\Lambda$ and is parallel to $\Gamma_{\rm wedge}$.
@@
    If state $(O)$ at $P_{1}$ is supersonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|>c_{O}$, the solution has the supersonic reflection structure as in Fig. 7 at point $P_{1}$. If state $(O)$ at $P_{1}$ is subsonic or sonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|\leq c_{O}$, the solution has the subsonic reflection structure as in Fig. 7.
@@
    The solution coincides with states $(\infty)$, $(O)$, and $(N)$ in their respective regions. Specifically,
@@
    and limξ∈Ω, ξ→P1(ρ, v)(ξ)=(ρO, vO)(P1)  if state (O) at $P_{1}$ is subsonic or sonic.
@@
We remark that, for the supersonic reflections, it follows from (ii)–(iv) of the definition above and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ is continuous across the sonic arcs $\Gamma_{\rm sonic}^{N}$ and $\Gamma_{\rm sonic}^{O}$.
@@
Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
@@
Then we can follow the proof of Theorem 4.10 to show that the regular solutions satisfy the assumptions of Theorem 3.2 similarly. The only (slight) difference is in showing that condition (b) of Lemma 4.1 holds. We need to show that $\Gamma_{\rm shock}$ is not a straight segment. Let us first consider the supersonic Prandtl reflection case; see Fig. 7. Arguing as in the proof of Theorem 4.10, we show that, at point $P_{4}$ (resp. $P_{5}$), curve $\Gamma_{\rm shock}$ is tangential to line $\Gamma_{\rm shock}^{N}$ (resp. $\Gamma_{\rm shock}^{O}$). If $\Gamma_{\rm shock}$ is a straight segment, we obtain that $\Gamma_{\rm shock}^{N}$ and $\Gamma_{\rm shock}^{O}$ lie in the same line. However, $\Gamma_{\rm shock}^{N}$ is parallel to $\Gamma_{\rm wedge}$ with a positive distance from it, while $\Gamma_{\rm shock}^{O}$ passes through point $P_{1}\in\Gamma_{\rm wedge}$. This contradiction shows that $\Gamma_{\rm shock}$ is not a straight segment. In the case of subsonic or sonic Prandtl reflection as in Fig. 7, we argue similarly, except that we consider point $P_{1}$ instead of point $P_{5}$ and use the fact that, in the subsonic or sonic case, $\Gamma_{\rm shock}$ is parallel to $\Gamma_{\rm shock}^{O}$ at $P_{1}$ by the last equality in condition (iv) of the definition of admissible solutions. Then we conclude in the same way as in the case of supersonic Prandtl reflection. Thus, we do not repeat the similar arguments for the proof. ∎
@@ 4.3. Lower Regularity of the Shock Diffraction Solutions of the Lighthill Problem
The third problem is the Lighthill problem, i.e., the shock diffraction problem, for the isentropic Euler system. As discussed in [16] for the potential flow and shown in Fig. 8, the Lighthill shock diffraction problem arises as a straight incident shock passes through a wedge stepping down.
Figure 8. Lighthill shock diffraction problem
Initially, we consider two piecewise constant Riemann data with the left state $(1)$: $(\rho_{1},u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$ and the right state $(0)$: $(\rho_{0},0,0)$, separated by a vertical shock, which hits the wedge at the wedge-corner $P_{1}$. Then the Lighthill shock diffraction problem evolves in a self-similar structure as time goes on. In the self-similar coordinates $\boldsymbol{\xi}$}, the incident shock $\Gamma_{\rm shock}^{1}$ is given by $\xi_{1}=\xi_{1}^{0}$. By a straightforward calculation, it follows from the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12) that the location of the incident shock satisfies $0<\xi_{1}^{0}<c_{1}$, where $c_{1}=\rho_{1}^{\frac{\gamma-1}{2}}$ is the sonic speed of state $(1)$. Thus, as shown in Fig. 8, the incident shock $\Gamma_{\rm shock}^{1}$ interacts with the sonic circle $\Gamma_{\rm sonic}$ of state $(1)$ and becomes a transonic shock $\Gamma_{\rm shock}$, and then $\Gamma_{\rm shock}$ meets wedge $\Gamma_{\rm wedge}^{0}$ perpendicularly.
Let $\Lambda_{0}$, $\Lambda_{1}$, and $\Omega$ be defined, as shown in Fig. 8. Let $\Lambda=\overline{\Lambda_{0}\cup\Lambda_{1}\cup\Omega}$. An entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in $\Lambda$ in the sense of Definition 2.2 is called a regular shock diffraction solution of the Lighthill problem for the isentropic Euler system if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
@@
    $\rho,{\mathbf{v}}\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})\setminus\{P_{3}\}\right)\cap C^{0,1}\left({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega}\right)$ for some $\sigma>0$;
@@
Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
@@ 4.4. Lower Regularity of the Riemann Solutions with Four-Shock Interactions
where $\theta$ is the polar angle of point ${\mathbf{x}}\in\mathbb{R}^{2}$ and the four parameters $\theta_{12},\theta_{32},\theta_{34},\theta_{14}\in(0,\frac{\pi}{2})$. On each $\Lambda_{i}$, suitable constant states $(i)$ with values $(\rho_{i},{\mathbf{v}}_{i})$ are given for the Riemann initial data such that any two neighboring states are connected by exactly one planar shock discontinuity, which satisfies the Rankine-Hugoniot conditions and the entropy condition for the isentropic Euler system. As introduced in [12], under the symmetry assumption that $\theta_{12}=\theta_{14}=\theta_{1}$, $\theta_{32}=\theta_{34}=\theta_{2}$, and states in $\Lambda_{2}$ and $\Lambda_{4}$ are the same constants, and under a structure assumption that one forward shock is generated between states $(1)$ and $(j)$ and one backward shock is generated between states $(3)$ and $(j)$, governed by the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12), we expect the Riemann problem develops a Riemann solution with the structure as shown in Fig. 9.
Figure 9. Riemann problem with four-shock interactions
On the symmetry line that is the dashed line shown in Fig. 9, the velocity satisfies (2.13) with $\boldsymbol{\nu}$ being the vertical direction. At the interaction points $P_{1}$ and $P_{4}$, based on an argument similar to the one for the regular shock reflection problem at the reflection point $P_{0}$, with the symmetry line and state $(2)$ in Fig. 9 corresponding to $\Gamma_{\rm wedge}$ and state $(1)$ in Fig. 3, there exists a detachment angle $\theta^{\rm d}\in(0,\frac{\pi}{2})$ depending on the data such that, for $\theta_{1},\theta_{2}\in(0,\theta_{\rm w}^{\rm d})$, there exist two constant states $(5)$ and $(6)$ in $\Lambda_{5}$ and $\Lambda_{6}$, respectively, satisfying the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12). One corresponds to the strong shock and the other to the weak shock. We always select the weak one because it expects to be stable in general. Then there exists a sonic angle $\theta^{\rm s}\in(0,\theta^{\rm d}]$ such that the state at $P_{1}$ (or $P_{4}$) is pseudo-supersonic if $\theta_{1}\in(0,\theta^{\rm s})$ (or $\theta_{2}\in(0,\theta^{\rm s})$) and is pseudo-subsonic if $\theta_{1}\in(\theta^{\rm s},\theta^{\rm d}]$ (or $\theta_{2}\in(\theta^{\rm s},\theta^{\rm d}]$).
@@
[…]
@@
Let $\Lambda_{i}$ for $i=1,\cdots,6$, if they exist, and let $\Omega$ be defined as shown in Fig. 9. Let $\Lambda=\mathbb{R}^{2}$. An entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in $\Lambda$ in the sense of Definition 2.2 is called a regular Riemann solution with four-shock interactions for the isentropic Euler system if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
@@
[…]
    where limξ∈Ω, ξ→P4(ρ, v)(ξ)=(ρ5, v5)(P4)  if state (5) at $P_{4}$ is subsonic or sonic, limξ∈Ω, ξ→P1(ρ, v)(ξ)=(ρ6, v6)(P1)  if state (6) at $P_{1}$ is subsonic or sonic.
@@
If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both supersonic, then $M=6$, ${N_{1}}=0$, and ${N_{2}}=4$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig. 9, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{6}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{4}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both subsonic, then $M=4$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,2,3,4$) being given as in Fig. 9, $\Lambda_{5}$ and $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{1}$.
If state $(5)$ at $P_{4}$ is supersonic and state $(6)$ at $P_{1}$ is subsonic, then $M=5$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,\cdots,5$) being given as in Fig. 9, $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{3}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{1}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
It state $(5)$ at $P_{4}$ is subsonic and state $(6)$ at $P_{1}$ is supersonic, then $M=5$, ${N_{1}}=0$ and ${N_{2}}=3$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig. 9, $\Lambda_{5}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$.
@@
    For some $\sigma>0$,
    $(\rho,{\mathbf{v}})\in$ $\,C^{1}\big(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})\setminus\{P_{2},P_{3},P_{5},P_{6}\}\big)$
    $\,\cap C^{0,1}\big({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega}\big),$
    where $\Gamma_{\rm sonic}^{i}=\emptyset$ when it does not exist for $i=1,2$;
@@
Then ${\mathbf{v}}\notin H^{1}(\Omega)$.